OpenStudy (anonymous):

what is the slope of a line parallel to this one with the points (0,3) and (3,2) plotted on a graph a)3/2 b)1/3 c)-1/2 d)-3

OpenStudy (anonymous):

What do we know about the slope of parallel lines?

OpenStudy (anonymous):

@domu isnt it y=mx+b

OpenStudy (anonymous):

That's the equation of a line. In there is the slope and y-intercept, which of the letters, y,m,x, or b, represents the slope?

OpenStudy (anonymous):

@domu is it m?

OpenStudy (anonymous):

Exactly! So if we have 2 parallel lines, what can we say about their slopes?

OpenStudy (anonymous):

that they are the same @domu

OpenStudy (anonymous):

yup yup, so now we just need to calculate the slope of the line passing through the two points given. How would we do that?

OpenStudy (anonymous):

hmmm how did you get there, can you show me your work?

OpenStudy (anonymous):

sorry i honestly have no idea what im doing lol hold on let me try it again

OpenStudy (anonymous):

@domu

OpenStudy (anonymous):

lol no problem

OpenStudy (anonymous):

i cant do this lol @domu is it -3 ??

OpenStudy (anonymous):

Given 2 points (x1,y1) and (x2,y2) how do we calculate the slope of the line passing through those 2 points? Is there a formula?

OpenStudy (anonymous):

@domu i got 1/3!

OpenStudy (anonymous):

@elizabethvilleda \[(y-y1/y2-y1)=(x-x1/x2-x1)\]

OpenStudy (anonymous):

from this we get the equation of this line

OpenStudy (anonymous):

isnt the slope=-1/3?

OpenStudy (texaschic101):

slope(m) = (y2 - y1) / (x2 - x1)

OpenStudy (anonymous):

lol thats what i meant sorry

OpenStudy (anonymous):

@elizabethvilleda -1/3?

OpenStudy (texaschic101):

(0,3) -- x1 = 0, y1 = 3 (3,2) -- x2 = 3, y2 = 2 m = (y2 -y1) / (x2 - x1) m = (2 - 3) / (3 - 0) m = -1/3. correct...slope is -1/3

OpenStudy (anonymous):

@texaschic101 yup it is